Mr.Belen is 30 years older than his son. How old are the father and the son if Mr. Belen is 10 more than thrice as old as his son.
Answer by mananth(16946)

son ----- x
belen = x+30
...
x+30 = 3x+10
3x-x=30-10
2x=20
/2
x=10 son's age
Belen's age = 10+30 =40 years